Experiment to find the enthalpy change for the reaction

Zn(s) + Cu2+(aq) → Cu(s) + Zn2+(aq)


By adding 25grams of zinc powder to a measured amount of aqueous copper(ll) sulphate, and measuring the temperature change over a period of time, you can then calculate the enthalpy change of the reaction.


1
Expert's answer
2021-03-22T06:27:59-0400

25g of Zn was added to 25cm³ of 1M CuSO4 solution in a polystyrene cup over 8 minutes. The temperature of the solution was observed to decrease from 68°C to 25°C.

Moles of Zn= 25/65.4= 0.38mol

Molar mass of CuSO4= 160g/mol

Mole of CuSO4= c x V = 1 x 25/1000= 0.025mol

Mass of CuSO4= mole x Molar mass= 0.025x160= 4g

q= mc∆T= 4 x 4.2 x (25-68)= -722.4J

This is the heat lost by CuSO4.

Since the temperature decreases, heat is negative and process is exothermic

∆H= q/n= -722.4/0.025= -28896J/mol